# The Role of Deductive and Inductive Reasoning in Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities in reasoning tasks, yet their reliance on static prompt structures and limited adaptability to complex scenarios remains a significant challenge. In this paper, we propose the Deductive and InDuctive(DID) method, a novel framework that enhances LLM reasoning by dynamically integrating both deductive and inductive reasoning approaches. Drawing from cognitive science principles, DID implements a dual-metric complexity evaluation system that combines Littlestone dimension and information entropy to precisely assess task difficulty and guide decomposition strategies. DID enables the model to progressively adapt its reasoning pathways based on problem complexity, mirroring human cognitive processes. We evaluate DID’s effectiveness across multiple benchmarks, including the AIW and MR-GSM8K, as well as our custom Holiday Puzzle dataset for temporal reasoning. Our results demonstrate significant improvements in reasoning quality and solution accuracy - achieving 70.3% accuracy on AIW (compared to 62.2% for Tree of Thought), while maintaining lower computational costs. The success of DID in improving LLM performance while preserving computational efficiency suggests promising directions for developing more cognitively aligned and capable language models. Our work contributes a theoretically grounded, input-centric approach to enhancing LLM reasoning capabilities, offering an efficient alternative to traditional output-exploration methods.
@@
Chengkun Cai1∗,  Xu Zhao1∗,  Haoliang Liu2∗,  Zhongyu Jiang3,  Tianfang Zhang4, Zongkai Wu5,  Jenq-Neng Hwang3,  Serge Belongie6,  Lei Li3,6† thanks: * Equal contribution. thanks: $\dagger$ Corresponding author: lilei@di.ku.dk 1University of Edinburgh  2University of Manchester  4Tsinghua University 5FancyTech  3University of Washington  6University of Copenhagen
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4, have transformed natural language processing by excelling in tasks such as language translation, summarization, and question-answering (OpenAI, 2023), particularly in reasoning tasks and few-shot learning. However, their reliability in problem-solving remains debatable. While Zhou et al. (2024) notes that scaling and fine-tuning can introduce unpredictable errors even in simple tasks, recent methodologies like Chain of Thought (CoT) (Wei et al., 2022b) have shown significant improvements in arithmetic and symbolic reasoning tasks (Li et al., 2024). Studies have demonstrated that LLMs can achieve high accuracy in multi-step reasoning when guided by structured approaches like CoT and self-consistency (Bubeck et al., 2023; Wang et al., 2022). Additionally, techniques such as reinforcement learning from human feedback (RLHF) have proven effective in reducing harmful or inaccurate outputs (Ouyang et al., 2022; Christiano et al., 2017).
Despite these advances, LLMs face significant challenges with complex and evolving tasks due to their reliance on static prompt structures and pre-learned patterns. This limitation manifests in tasks requiring logical reasoning, such as calculating family relationships or performing numerical comparisons (Nezhurina et al., 2024). Unlike human problem-solving, which dynamically adjusts strategies based on task complexity through inductive and deductive reasoning (Sloman, 2009), LLMs often struggle to adapt their reasoning processes to novel situations (Marcus, 2020; Hendrycks et al., 2020).
This adaptability gap becomes particularly evident in tasks requiring dynamic adjustment or incremental problem-solving. While existing approaches like CoT (Wei et al., 2022b), Tree-of-Thought (ToT) (Yao et al., 2024), Temperature-Tree-of-Thought (T2oT) (Cai et al., 2024), and Graph-of-Thought (GoT) (Besta et al., 2024) have made progress through extensive output exploration, they often incur significant computational costs. For instance, ToT achieves 62.2% accuracy on the AIW benchmark but requires substantial output token generation for exploring multiple reasoning paths, resulting in higher computational overhead (\$0.0038 per case compared to \$0.0022 for CoT).
@@
We validate DID’s effectiveness on established benchmarks including AIW and MR-GSM8K (Wei et al., 2022a; Cobbe et al., 2021), as well as our custom Holiday Puzzle dataset focusing on holiday date calculations. As shown in Figure 1, our empirical results demonstrate significant improvements in both solution accuracy and reasoning quality, achieving 70.3% accuracy on AIW (compared to 62.2% for ToT) while maintaining lower computational costs (\$0.0031 vs \$0.0038 per case). This work makes the following key contributions:
@@ 2 Related Works
While LLMs like GPT-4 excel at tasks such as text generation, they struggle with logical reasoning and complex deduction (OpenAI, 2023; Nezhurina et al., 2024). Techniques like CoT (Wei et al., 2022b), ToT (Yao et al., 2024), and GoT (Besta et al., 2024) improve reasoning by structuring problems, but they require extensive prompt engineering and lack real-time adaptability. The DID framework addresses these limitations by dynamically integrating inductive and deductive reasoning, improving adaptability and consistency in complex tasks (Marcus, 2020; Gershman et al., 2015).
Recent advances in LLM reasoning capabilities have explored different approaches to enhancing model performance at test time. Models like DeepSeek-R1 Guo et al. (2025) and o1-ioi El-Kishky et al. (2025) achieve impressive results by leveraging reinforcement learning during pre-training and extending reasoning paths during inference - with DeepSeek-R1 reaching 79.8% accuracy on AIME 2024 through naturally emerged reasoning behaviors, and o1-ioi employing sophisticated test-time compute strategies to evaluate multiple solution candidates. Our DID framework takes a complementary approach by explicitly structuring input complexity during inference. This framework maintains explicit control through structured decomposition without requiring extensive computational resources or complex test-time evaluation strategies, potentially offering a more efficient path to enhanced reasoning capabilities while remaining easily accessible to the broader research community.
@@
Recent work links inductive inference to online learning theory. (Lu, 2024) demonstrate that inductive inference is possible for hypothesis classes decomposable into countable unions with finite Littlestone dimension. This result extends classical induction models, such as Solomonoff’s (Solomonoff, 1964). The connection between Littlestone dimension and learning complexity informs the DID framework, suggesting that decomposing tasks into simpler components can enhance learning and generalization.
@@ 3.1.1 Littlestone Dimension and Beyond
Both problems share the same Littlestone dimension as they follow identical reasoning patterns. Yet, LLMs consistently perform better on Problem A. This discrepancy arises from several theoretical considerations:
@@ 3.1.2 Information Entropy Component
where $|x_{i}|$ represents the absolute value of each numerical variable in the problem. This formulation:
@@ 3.1.3 Problem Complexity Evaluation
The decomposition process starts by creating a base case with reduced dimension ($d-2$), achieved by setting certain variables to zero. This simplification maintains the essential reasoning structure while reducing the problem’s complexity to its most basic form. From this foundation, the algorithm iteratively constructs $N=\lceil C(p)\rceil$ subproblems of increasing complexity, where $C(p)$ represents the overall problem complexity.
@@ 3.2 De-In-Ductive (DID) Framework
Figure 2: Comparison of reasoning approaches in Large Language Models (LLMs) including the IO method, Chain of Thought (CoT) prompting, Tree of Thought (ToT) prompting, and the De-In-Ductive (DID) framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
@@
  • Hypothesis Generation: Through progressive exposure to increasingly complex examples, the model generates and refines hypotheses about underlying problem structures. Each subproblem serves as a training instance for pattern recognition.
@@
  • Hierarchical Problem Solving: The deductive process follows the complexity hierarchy established during induction, ensuring that solutions build systematically on previously verified patterns.
The results from deductive applications inform and refine the inductive pattern recognition process, creating a continuous learning cycle that enhances the model’s problem-solving capabilities.
@@
The De-In-Ductive (DID) method seamlessly integrates with various LLM architectures and existing techniques such as CoT prompting (Wei et al., 2022b). Through its structured framework combining inductive and deductive reasoning, DID enhances these methods by providing dynamic reasoning strategies and guided incremental reasoning, while maintaining computational efficiency. This approach creates a more flexible framework for LLMs to address complex problems without significant overhead.
@@ 4.2 Tasks and Results
Figure 3: Comparison of reasoning approaches in Large Language Models (LLMs) including the IO method, Chain of Thought (CoT) prompting, and the De-In-Ductive (DID) framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
The AIW dataset focuses on evaluating logical reasoning and deduction abilities through family relationship problems. We manually curated 113 unique problems after removing duplicates and existing prompts, with results averaging over 20 runs. In this task, DID demonstrates consistent superiority across all models:
@@
MR-GSM8K extends the GSM8K benchmark with meta-reasoning tasks, requiring models to identify and explain errors in provided solutions. Results show consistent performance across models:
@@
Figure 4: Comparison of reasoning approaches in Large Language Models (LLMs) including the IO method, Chain of Thought (CoT) prompting, and the De-In-Ductive (DID) framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
@@
This custom dataset comprises 20 holiday arrangement problems, testing models’ ability to calculate actual holiday days while accounting for weekends and compensatory workdays. Results demonstrate:
@@ 5 Conclusion
In this work, we introduced the De-In-Ductive (DID) method, a novel framework that dynamically integrates inductive and deductive reasoning to enhance the adaptability and reasoning capabilities of Large Language Models (LLMs). By leveraging cognitive science principles, the DID framework allows LLMs to evolve their problem-solving strategies in response to task complexity, overcoming the rigidity of static prompt structures. Through extensive empirical validation on both standard benchmarks and our custom Holiday Puzzle dataset, we demonstrated significant improvements in accuracy and reasoning quality, achieved without excessive computational costs. The success of DID in improving LLM reasoning while maintaining computational efficiency suggests promising directions for future research in making language models more cognitively aligned and capable of sophisticated reasoning.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities in reasoning tasks, yet their reliance on static prompt structures and limited adaptability to complex scenarios remains a major challenge. In this paper, we propose the Deductive and InDuctive(DID) method, a novel framework that enhances LLM reasoning by dynamically integrating both deductive and inductive reasoning approaches. Drawing from cognitive science principles, DID implements a dual-metric complexity evaluation system that combines Littlestone dimension and information entropy to precisely assess task difficulty and guide decomposition strategies. DID enables the model to progressively adapt its reasoning pathways based on problem complexity, mirroring human cognitive processes. We evaluate DID’s effectiveness across multiple benchmarks, including the AIW, MR-GSM8K, and our custom Holiday Puzzle dataset for temporal reasoning. Our results demonstrate great improvements in reasoning quality and solution accuracy - achieving 70.3% accuracy on AIW (compared to 62.2% for Tree of Thought), while maintaining lower computational costs.
@@
Chengkun Cai1thanks: Equal contribution.  Xu Zhao [^1] footnotemark: 1  Haoliang Liu211footnotemark: 1  Zhongyu Jiang3 Tianfang Zhang4  Zongkai Wu5thanks: Project Leader.  Jenq-Neng Hwang3  Lei Li3,6thanks: Corresponding author: lilei@di.ku.dk 1University of Edinburgh  2University of Manchester  3University of Washington 4Tsinghua University  5Skai Intelligence  6University of Copenhagen
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4, have transformed natural language processing by excelling in tasks such as language translation, summarization, and question-answering (OpenAI, 2023), particularly in reasoning tasks and few-shot learning. However, their reliability in problem-solving remains debatable. While Zhou et al. (2024) notes that scaling and fine-tuning can introduce unpredictable errors even in simple tasks, recent methodologies like Chain of Thought (CoT) (Wei et al., 2022) have shown substantial improvements in arithmetic and symbolic reasoning tasks (Li et al., 2024). Studies have demonstrated that LLMs can achieve high accuracy in multi-step reasoning when guided by structured approaches like CoT and self-consistency (Bubeck et al., 2023; Wang et al., 2022). Additionally, techniques such as reinforcement learning from human feedback (RLHF) have proven effective in reducing harmful or inaccurate outputs (Ouyang et al., 2022; Christiano et al., 2017).
Despite these advances, LLMs face substantial challenges with complex and evolving tasks due to their reliance on static prompt structures and pre-learned patterns. This limitation manifests in tasks requiring logical reasoning, such as calculating family relationships or performing numerical comparisons (Nezhurina et al., 2024). Unlike human problem-solving, which dynamically adjusts strategies based on task complexity through inductive and deductive reasoning (Sloman, 2009), LLMs often struggle to adapt their reasoning processes to novel situations (Marcus, 2020; Hendrycks et al., 2020).
This adaptability gap becomes particularly evident in tasks requiring dynamic adjustment or incremental problem-solving. While existing approaches like CoT (Wei et al., 2022), Tree-of-Thought (ToT) (Yao et al., 2024), Temperature-Tree-of-Thought (T2oT) (Cai et al., 2024), and Graph-of-Thought (GoT) (Besta et al., 2024) have made progress through extensive output exploration, they often incur considerable computational costs. For instance, ToT achieves 62.2% accuracy on the AIW benchmark but requires substantial output token generation for exploring multiple reasoning paths, resulting in higher computational overhead (\$0.0038 per case compared to \$0.0022 for CoT).
@@
We validate DID’s effectiveness on established benchmarks including AIW and MR-GSM8K (Nezhurina et al., 2024; Zeng et al., 2023), as well as our custom Holiday Puzzle dataset focusing on holiday date calculations. As shown in Figure 1, our empirical results demonstrate notable improvements in both solution accuracy and reasoning quality, achieving 70.3% accuracy on AIW (compared to 62.2% for ToT) while maintaining lower computational costs (\$0.0031 vs \$0.0038 per case). This work makes the following key contributions:
@@ 2 Related Works
While LLMs like GPT-4 excel at tasks such as text generation, they struggle with logical reasoning and complex deduction (OpenAI, 2023; Nezhurina et al., 2024). Techniques like CoT (Wei et al., 2022), ToT (Yao et al., 2024), and GoT (Besta et al., 2024) improve reasoning by structuring problems, but they require extensive prompt engineering and lack real-time adaptability. The DID framework addresses these limitations by dynamically integrating inductive and deductive reasoning, improving adaptability and consistency in complex tasks (Marcus, 2020; Gershman et al., 2015).
Recent comprehensive surveys by Giadikiaroglou et al. (2024) and Liu et al. (2023) provide thorough analyses of LLM reasoning capabilities in puzzle-solving and mathematical domains, highlighting both progress and persistent challenges in structured reasoning tasks. Recent advances in LLM reasoning capabilities have explored different approaches to enhancing model performance at test time. Models like DeepSeek-R1 (Guo et al., 2025) and o1-ioi (El-Kishky et al., 2025) achieve impressive results by leveraging reinforcement learning during pre-training and extending reasoning paths during inference - with DeepSeek-R1 reaching 79.8% accuracy on AIME 2024 through naturally emerged reasoning behaviors, and o1-ioi employing sophisticated test-time compute strategies to evaluate multiple solution candidates.
Recent advancements in improving LLM reasoning have explored diverse strategies. While multi-agent frameworks like Jin et al. (2025) demonstrate superior performance through collaborative exploration of reasoning paths, they incur increased computational overhead due to multiple model calls. Other approaches, such as "learning from teaching regularization" (Jin et al., 2024) and Self-Explore (Hwang et al., 2024), enhance reasoning by incorporating structured examples or fine-grained rewards during training. However, these methods necessitate model fine-tuning or multiple model instances, whereas our DID framework distinguishes itself by improving reasoning through structured prompting of a single model instance without additional training.
@@
Recent work links inductive inference to online learning theory. Lu (2024) demonstrate that inductive inference is possible for hypothesis classes decomposable into countable unions with finite Littlestone dimension. This result extends classical induction models, such as Solomonoff’s (Solomonoff, 1964). The connection between Littlestone dimension and learning complexity informs the DID framework, suggesting that decomposing tasks into simpler components can enhance learning and generalization.
@@ 3.1.1 Littlestone Dimension and Beyond
Both problems share the same Littlestone dimension, as they follow identical reasoning patterns. However, LLMs consistently perform better on Example 1. This discrepancy arises from several theoretical considerations:
@@ 3.1.2 Information Entropy Component
where $|x_{i}|$ represents the absolute value of each numerical variable in the problem. Importantly, only variables that are directly relevant to solving the problem are included in this calculation. For instance, in the problem "Alice has 3 brothers and 6 sisters. How many sisters does Alice’s brother have?", only the number of brothers (3) and sisters (6) would be considered in the entropy calculation. This formulation:
@@ 3.1.3 Problem Complexity Evaluation
The decomposition process starts by creating a base case with reduced dimension ($d-2$), achieved by setting certain variables to zero. Specifically, this means eliminating key decision points in the reasoning chain by simplifying the problem structure, for example, changing "Alice has 3 brothers and 6 sisters" to "Alice has 0 sisters and 1 brother" to reduce the problem’s Littlestone dimension. This simplification maintains the essential reasoning structure while reducing the problem’s complexity to its most basic form. From this foundation, the algorithm iteratively constructs $N=\lceil C(p)\rceil$ subproblems of increasing complexity, where $C(p)$ represents the overall problem complexity.
@@ 3.2 De-In-Ductive (DID) Framework
Figure 2: Comparison of reasoning approaches in LLMs including the IO method, Chain of Thought (CoT) prompting, Tree of Thought (ToT) prompting, and the De-In-Ductive (DID) framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
@@
  • Hypothesis Generation: Through progressive exposure to increasingly complex examples, the model generates and refines hypotheses about the underlying structure of the problem. Each subproblem serves as a training instance for pattern recognition.
@@
  • Hierarchical Problem Solving: The deductive process follows the complexity hierarchy established during induction, ensuring that solutions are built systematically on previously verified patterns.
The results from deductive applications inform and refine the inductive pattern recognition process, creating a continuous learning cycle that enhances the model’s problem-solving capabilities. A complete step-by-step example of the DID framework is provided in Appendix B.
@@
The DID method seamlessly integrates with various LLM architectures and existing techniques such as CoT prompting (Wei et al., 2022). Through its structured framework combining inductive and deductive reasoning, DID enhances these methods by providing dynamic reasoning strategies and guided incremental reasoning, while maintaining computational efficiency. This approach creates a more flexible framework for LLMs to address complex problems without notable overhead.
@@ 4.2 Tasks and Results
Figure 3: Comparison of reasoning approaches in LLMs including the IO method, CoT prompting, and the DID framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
Figure 4: Comparison of reasoning approaches in LLMs including the IO method, CoT prompting, and the DID framework, highlighting the progression from direct output generation to dynamic inductive and deductive reasoning for more adaptive problem-solving.
The AIW dataset focuses on evaluating logical reasoning and deduction abilities through family relationship problems (Nezhurina et al., 2024). We manually curated 113 unique problems after removing duplicates and existing prompts, with results averaging over 20 runs. In this task, DID demonstrates consistent superiority across all models:
@@
MR-GSM8K extends the GSM8K benchmark with meta-reasoning tasks (Zeng et al., 2023), requiring models to identify and explain errors in provided solutions. Results show consistent performance across models:
@@
This custom dataset comprises 20 holiday arrangement problems, testing models’ ability to calculate actual holiday days while accounting for weekends and compensatory workdays. Detailed information about the dataset construction and representative examples are provided in Appendix A. Results demonstrate:
@@ 5 Conclusion
In this work, we introduced the De-In-Ductive (DID) method, a novel framework that dynamically integrates inductive and deductive reasoning to enhance the adaptability and reasoning capabilities of LLMs. By leveraging cognitive science principles, the DID framework allows LLMs to evolve their problem-solving strategies in response to task complexity, overcoming the rigidity of static prompt structures. Through extensive empirical validation on both standard benchmarks and our custom Holiday Puzzle dataset, we demonstrated substantial improvements in accuracy and reasoning quality, achieved without excessive computational costs. The success of DID in improving LLM reasoning while maintaining computational efficiency suggests promising directions for future research in making language models more cognitively aligned and capable of sophisticated reasoning.
@@ 6 Limitations
7 Acknowledgements
This work was supported in part by the Pioneer Centre for AI, DNRF grant number P1.
Appendix A Holiday Puzzle Dataset Details
The Holiday Puzzle dataset was created based on holiday arrangements in China over the past 10 years, specifically focusing on how special holidays (National Day, Spring Festival, Labor Day, Mid-Autumn Festival, etc.) are rescheduled. In China, the government employs a unique "work day adjustment" system to create longer consecutive holiday periods by rearranging working days and weekends. This practice often involves designating certain weekends as working days while extending official holidays, creating complex patterns where regular weekends are shifted, and compensatory workdays are inserted before or after holidays. This arrangement, while allowing for longer holiday periods, makes it challenging to calculate the actual number of holiday days versus regular weekend days.
A.1 Representative Examples
Prompt: This is a holiday arrangement from April 23, 2022 (Saturday) to May 15: rest 1, work 6, rest 5, work 3, rest 1, work 5, rest 2. Please tell me how many days I have rested because of the holiday, except for the weekend I was supposed to rest.
Right Answer: 1
Prompt: This is a holiday arrangement from January 1, 2022 (Saturday) to February 8: rest 3, work 4, rest 2, work 5, rest 2, work 5, rest 2, work 7, rest 7, rest 2. Please tell me how many days I have rested because of the holiday, except for the weekend I was supposed to rest.
Right Answer: 4
Appendix B Detailed DID Framework Example
We provide a complete example of the DID framework in action:
Problem: "Alice has 3 brothers and 6 sisters. How many sisters does Alice’s brother have?"
Step 1: Complexity Evaluation The LLM first analyzes the problem structure and identifies:
  • Littlestone dimension ($d$): 3 (requiring three key inferential steps)
  • Information entropy: $H(p)=\log_{2}((1+3)(1+6))=\log_{2}(28)\approx 4.8$
  • Overall complexity: $C(p)=d\cdot H(p)=3\cdot 4.8\approx 14.4$
Step 2: Problem Decomposition Following Algorithm 1, the problem is decomposed into a sequence of progressively complex subproblems:
Inductive Reasoning Phase:
  • Example 1 (d-1 Dimension): "Let’s start with the simplest version of the problem: Alice has 0 sisters and 1 brother. In this case, Alice’s brother has 1 sister (Alice)."
  • Example 2 (d-1 Dimension): "Now, let’s consider a slightly more complex scenario: Alice has 1 sister and 1 brother. In this case, Alice’s brother has 2 sisters (Alice and her sister)."
  • Example 3 (d Dimension): "Let’s examine one more scenario: Alice has 2 sisters and 1 brother. In this case, Alice’s brother has 3 sisters (Alice and her two sisters)."
  • Pattern Recognition: "From these examples, we can deduce a general rule: The number of sisters Alice’s brother has = X + 1, where X = the number of sisters Alice has. This rule holds true regardless of the number of brothers Alice has (Y), because we’re only concerned about the number of sisters from the brother’s perspective."
Step 3: Deductive Reasoning "Now I’ll apply this formula to our original problem: Alice has 6 sisters (X = 6). Therefore, Alice’s brother has X + 1 = 6 + 1 = 7 sisters."
Step 4: Solution Verification "To verify: Alice’s brother has the same siblings as Alice, except himself.
  • Alice has 3 brothers and 6 sisters
  • From her brother’s perspective, he has 2 brothers (the other brothers, excluding himself) and 7 sisters (the 6 original sisters plus Alice)
Therefore, Alice’s brother has 7 sisters."
Final Answer: 7